# A Chinese Continuous Sign Language Dataset Based on Complex Environments

### v1

@@ IV-C Experimental results
In Table VII, we compare the performance of various CSLR methods on the CE-CSL dataset. As shown in Table VII, the test results of our proposed model on CE-CSL indicate its superior performance. We also analyze the performance of TFNet on CE-CSL from both quantitative and qualitative perspectives.
Quantitative Analysis: To evaluate the effectiveness of the proposed model, we conduct comparative experiments on the CE-CSL dataset. The results reveale significant advantages of our method. According to the data in Table VII, our model achieve WER of 42.1% and 41.9% on the validation and test sets respectively. Compared to the closest CSLR model MAM-FSD, which have WER values of 44.9% and 44.7% on the validation and test sets, our model demonstrates a reduction of 2.8%, highlighting the superiority of our proposed model in sign language recognition.
Qualitative Analysis: To delve deeper and provide an intuitive demonstration of the performance of different models in recognizing Chinese continuous sign language datasets in complex environments, we select several CSLR models and analyze their test results on the CE-CSL dataset. The analyzed results in Table VIII, aiming to offer a more intuitive perspective. With the results in Table VIII, we find that the TFNet model has the lowest recognition error rate, marking it as having the best recognition performance. This result reflects the model’s excellent capability in handling complex backgrounds and continuous sign language videos.
@@ IV-D Ablation experiment
Ablation of different domain sequence features. We conduct experiments by combining temporal, spectral, and time-frequency domain sequence features after frame-level feature extraction to investigate the role of different domain sequence feature combinations in CSLR. The experimental results are shown in Table IX. As can be observed from Table IX, when using only temporal domain sequence features for recognition, the WERs for the validation and test sets are 19.5% and 19.1%, respectively. When using only spectral domain sequence features for recognition, the WERs are 19.7% and 20.0%, respectively. This indicates that using only temporal domain sequence features yields slightly better performance than using only spectral domain sequence features, with WERs reduced by 0.2% and 0.9%, respectively. Finally, by integrating temporal and spectral domain sequence features to form time-frequency domain sequence features for recognition, the results showed further improvements. Compared to using only temporal domain sequence features, the WER decreased by 0.8% and 0.5%, respectively, achieving even better results. This suggests that the integration of time-frequency domain sequence features can effectively enhance the performance of CSLR systems.
@@
Ablation of different sequence feature extractors. Under the condition that the frame-level feature extractors of different models remain unchanged, we investigate the effect of replacing their respective sequence feature extractors with time-frequency sequence feature extractors. The experimental results are shown in Table X. From Table X, it can be seen that for different models, recognition using time-frequency sequence features results in a decrease in WER on the test set compared to the original sequence features. Specifically, the WER of MAM-FSD is reduced by 0.2%, the WER of CorrNet is reduced by 0.3%, and the WER of SEN is reduced by 1.7%. On the validation set, only the BER of the CorrNet model increases by 0.4%, while the performance of the remaining models on the validation set improves with the use of the time-frequency sequence feature extractor. This experimental result proves the sophistication of the time-frequency sequence feature extractor we used.
@@
Ablation of Different Loss Combinations. We study the effects of various loss function combinations in the TFNet model by experimenting with different combinations of the CTC loss function $L_{CTC}$, the temporal domain sequence feature VAE loss function $L_{VAE_{t}}$ and the spectral domain sequence feature VAE loss function $L_{VAE_{f}}$. The experimental results are shown in Table XI. As can be observed from Table XI, using only $L_{CTC}$ results in WERs of 21.3% and 21.4% on the validation and test sets, respectively. When $L_{CTC}$ is combined with either of the VAE loss functions, there is an improvement in performance. Combining $L_{CTC}$ with $L_{VAE_{t}}$, the WERs on the validation and test sets decrease by 2.0% and 2.4%, respectively. Combining $L_{CTC}$ with $L_{VAE_{f}}$, the WERs on the validation and test sets decrease by 2.3% and 2.1%, respectively. When all three loss functions are combined, the WER reaches its minimum on both the validation and test sets, at 18.7% and 18.6%, respectively. The result indicates that combining the CTC loss function with the time-frequency domain sequence feature VAE loss functions can significantly enhance the performance of CSLR systems.
@@ V Conclusion
In this paper, we introduce a novel CE-CSL dataset designed for complex environments, aiming to facilitate the transition of CSLR from laboratory settings to real-life scenarios. The CE-CSL encompasses a rich variety of daily communication contexts, comprising a total of 5,988 continuous sign language videos, meticulously divided into a training set of 4,973 sign language videos, a validation set of 515 sign language videos, and a test set of 500 sign language videos. The dataset covers 3,515 Chinese words with a variety of video backgrounds, introducing more than 70 different environments, providing unprecedented complexity and realism for model training. Considering the impact of complex backgrounds on CSLR, we also propose the TFNet to achieve efficient and accurate recognition. On the validation and test sets of the CE-CSL, the TFNet achieves WERs of 42.1% and 41.9%, respectively. Experimental results indicate that the TFNet significantly improves recognition accuracy. Additionally, we applied our proposed method to three publicly available continuous sign language datasets and achieved highly competitive results.

### v2

@@ IV-C Experimental results
In Table VI, we compare the performance of various CSLR methods on the CE-CNSL dataset. As shown in Table VI, the test results of our proposed model on CE-CNSL indicate its superior performance. We also analyze the performance of TFNet on CE-CNSL from both quantitative and qualitative perspectives.
Quantitative Analysis: To evaluate the effectiveness of the proposed model, we conduct comparative experiments on the CE-CNSL dataset. The results reveal significant advantages of our method. According to the data in Table VI, our model achieve WER of 42.1% and 41.9% on the validation and test sets respectively. Compared to the closest CSLR model MAM-FSD, which have WER values of 44.9% and 44.7% on the validation and test sets, our model demonstrates a reduction of 2.8%, highlighting the superiority of our proposed model in sign language recognition.
Qualitative Analysis: To delve deeper and provide an intuitive demonstration of the performance of different models in recognizing Chinese continuous sign language datasets in complex environments, we select several CSLR models and analyze their test results on the CE-CNSL dataset. The analyzed results in Table VII, aiming to offer a more intuitive perspective. With the results in Table VII, we find that the TFNet model has the lowest recognition error rate, marking it as having the best recognition performance. This result reflects the model’s excellent capability in handling complex backgrounds and continuous sign language videos.
@@ IV-D Ablation experiment
Ablation of different domain sequence features. We conduct experiments by combining temporal, spectral, and time-frequency domain sequence features after frame-level feature extraction to investigate the role of different domain sequence feature combinations in CSLR. The experimental results are shown in Table VIII. As can be observed from Table VIII, when using only temporal domain sequence features for recognition, the WERs for the validation and test sets are 19.5% and 19.1%, respectively. When using only spectral domain sequence features for recognition, the WERs are 19.7% and 20.0%, respectively. This indicates that using only temporal domain sequence features yields slightly better performance than using only spectral domain sequence features, with WERs reduced by 0.2% and 0.9%, respectively. Finally, by integrating temporal and spectral domain sequence features to form time-frequency domain sequence features for recognition, the results showed further improvements. Compared to using only temporal domain sequence features, the WER decreased by 0.8% and 0.5%, respectively, achieving even better results. This suggests that the integration of time-frequency domain sequence features can effectively enhance the performance of CSLR systems.
@@
Ablation of different sequence feature extractors. Under the condition that the frame-level feature extractors of different models remain unchanged, we investigate the effect of replacing their respective sequence feature extractors with time-frequency sequence feature extractors. The experimental results are shown in Table IX. From Table IX, it can be seen that for different models, recognition using time-frequency sequence features results in a decrease in WER on the test set compared to the original sequence features. Specifically, the WER of MAM-FSD is reduced by 0.2%, the WER of CorrNet is reduced by 0.3%, and the WER of SEN is reduced by 1.7%. On the validation set, only the BER of the CorrNet model increases by 0.4%, while the performance of the remaining models on the validation set improves with the use of the time-frequency sequence feature extractor. This experimental result proves the sophistication of the time-frequency sequence feature extractor we used.
@@
Ablation of Different Loss Combinations. We study the effects of various loss function combinations in the TFNet model by experimenting with different combinations of the CTC loss function $L_{CTC}$, the temporal domain sequence feature VAE loss function $L_{VAE_{t}}$ and the spectral domain sequence feature VAE loss function $L_{VAE_{f}}$. The experimental results are shown in Table X. As can be observed from Table X, using only $L_{CTC}$ results in WERs of 21.3% and 21.4% on the validation and test sets, respectively. When $L_{CTC}$ is combined with either of the VAE loss functions, there is an improvement in performance. Combining $L_{CTC}$ with $L_{VAE_{t}}$, the WERs on the validation and test sets decrease by 2.0% and 2.4%, respectively. Combining $L_{CTC}$ with $L_{VAE_{f}}$, the WERs on the validation and test sets decrease by 2.3% and 2.1%, respectively. When all three loss functions are combined, the WER reaches its minimum on both the validation and test sets, at 18.7% and 18.6%, respectively. The result indicates that combining the CTC loss function with the time-frequency domain sequence feature VAE loss functions can significantly enhance the performance of CSLR systems.
@@ V Conclusion
In this paper, we introduce a novel CE-CNSL dataset designed for complex environments, aiming to facilitate the transition of CSLR from laboratory settings to real-life scenarios. The CE-CNSL encompasses a rich variety of daily communication contexts, comprising a total of 5,988 continuous sign language videos, meticulously divided into a training set of 4,973 sign language videos, a validation set of 515 sign language videos, and a test set of 500 sign language videos. The dataset covers 3,515 Chinese words with a variety of video backgrounds, introducing more than 70 different environments, providing unprecedented complexity and realism for model training. Considering the impact of complex backgrounds on CSLR, we also propose the TFNet to achieve efficient and accurate recognition. On the validation and test sets of the CE-CNSL, the TFNet achieves WERs of 42.1% and 41.9%, respectively. Experimental results indicate that the TFNet significantly improves recognition accuracy. Additionally, we applied our proposed method to three publicly available continuous sign language datasets and achieved highly competitive results.